# Coverage-based Fairness in Multi-document Summarization

### v1

@@ Abstract
Fairness in multi-document summarization (MDS) measures whether a system can generate a summary fairly representing information from documents with different social attribute values. Fairness in MDS is crucial since a fair summary can offer readers a comprehensive view. Previous works focus on quantifying summary-level fairness using Proportional Representation, a fairness measure based on Statistical Parity. However, Proportional Representation does not consider redundancy in input documents and overlooks corpus-level unfairness. In this work, we propose a new summary-level fairness measure, Equal Coverage, which is based on coverage of documents with different social attribute values and considers the redundancy within documents. To detect the corpus-level unfairness, we propose a new corpus-level measure, Coverage Parity. Our human evaluations show that our measures align more with our definition of fairness. Using our measures, we evaluate the fairness of thirteen different LLMs. We find that Claude3-sonnet is the fairest among all evaluated LLMs. We also find that almost all LLMs overrepresent different social attribute values.
@@ 3.2 Equal Coverage
However, two issues arise with summaries’ complex sentence structures. First, a summary sentence can combine information from several documents, making it difficult to attribute the sentence to any single document. Second, sentence lengths vary significantly based on social attribute values. The length difference can skew the coverage probability for different social attribute values $p(d,s|a=k)$. To address these issues, Equal Coverage splits the summary sentences $s\in S$ into multiple simpler sentences by prompting GPT-3.5 motivated by Bhaskar et al. (2023); Min et al. (2023). For example, compound sentences are split into simple sentences that describe a single fact. We denote the $j$-th summary sentence after split as $s_{j}$. Further details are in App. A.1.
Equal Coverage estimates the coverage probability for different social attribute values $p(d,s|a=k)$ and for a document $p(d,s)$ based on the probability $p(d_{i},s_{j})$ that a document $d_{i}$ is covered by a summary sentence $s_{j}$. The probability $p(d_{j},s_{k})$ is estimated as the entailment probability that the document $d_{j}$ entails the summary sentence $s_{k}$ by a textual entailment model Williams et al. (2018). Motivated by Laban et al. (2022), Equal Coverage divides documents into chunks of $M$ words. The $l$-th chunk of the document $d_{j}$ is denoted as $d_{j,l}$. The entailment model estimates the probability $p(d_{j},s_{k})$ as the maximum entailment probability $p(d_{j,l},s_{k})$ between the chunk $d_{j,l}$ and the summary sentence $s_{k}$:
@@ 4.2 Implementation Details
To measure the difficulty of obtaining fair summaries on different datasets, we estimate lower bounds, Lowergre, and upper bounds, Uppergre, for both fairness measures by greedily extracting sentences that minimize or maximize the measures. Please note that Lowergre and Uppergre are not baselines but serve as empirical bounds for our measures. The details of these bounds are in App.A.9.
@@ 5.2 Summary-level Fairness Evaluation
Previous works by Zhang et al. (2023) and Lei et al. (2024) evaluate the fairness of LLMs using Proportional Representation. When evaluating the fairness on sentiments in the review domain, our results using Equal Coverage are consistent with these works in the finding that GPT-4 is fairer than GPT-3.5, and both are fairer than COOP. However, they find that Llama2-13b is the fairest, while our results show that Llama2-13b and Llama2-70b are comparably fair. When evaluating the fairness on political ideologies in the tweet domain, our results are consistent with these works on that smaller models are fairer for Llama2. However, they find that GPT-4 is fairer than GPT-3.5 while we find that GPT-3.5 is fairer than GPT-4. We also show that GPT-4 is fairer than Llama2-7b, which contrasts the previous works. As shown in Sec. 1, Equal Coverage considers redundancy common in MDS, suggesting better reflects fairness in MDS.
@@ 5.4 Fainess under Different Distributions of Social Attributes
We perform experiments to evaluate whether the fairness of LLMs changes under different distributions of social attribute values in input document sets. For this, we divide all samples $G$ into $K$ non-overlapping sets: $\{G_{1},...G_{K}\}$ based on distributions of social attribute values. Each set $G_{k}$ includes samples where most documents $d\in D$ have a social attribute value of $k$. We denote the set $G_{k}$ as the set dominated by social attribute value $k$. We show results for differences of the summary-level fairness in App. A.10. To measure differences of the corpus-level fairness measured by Coverage Parity under different distributions, we use maximum differences of average coverage probability differences for the same social attribute value $\mathbb{E}(C_{j})$ (Sec. 3.3) on different sets $G_{k}$. The results are in Table 4.
@@ A.1 Split and Rephrase Summary Sentences
Motivated by Bhaskar et al. (2023); Min et al. (2023), we split all summary sentences into simple sentences. The goal of this step is to ensure that each summary sentence after the split only discusses a single fact. Specifically, compound sentences are split into simple sentences, while sentences with compound subjects or objects are split into sentences with simple subjects or objects. After the split, summary sentences are then rephrased to remove the reported speech, like ‘documents say what’. For splitting and rephrasing, we prompt GPT-3.5 with demonstrations. We show the example prompt and the example summary sentences after splitting and rephrasing in Fig. 2.
@@ A.4 Datasets
Mohammad et al. (2016) consists of tweets with labels of stance toward a short phrase such as Climate Change or Hillary Clinton. We cluster tweets about the same short phrase based on their TFIDF similarity into clusters. We then divide these clusters into input document sets of 20 tweets about the same short phrase.
@@ A.7 Fair Summarization Prompts
A.8 Correlation between Equal Coverage and Proportional Representation
Amazon MITweet Article Bias SemEval News Stance
GPT-3.5 -0.025 -0.137 0.042 0.097 0.046
Llama2-70b 0.015 -0.091 0.172 0.078 -0.011
Mixtral-8x7b -0.006 -0.085 0.320 -0.072 0.083
Claude3-haiku -0.066 -0.026 0.071 -0.006 0.063
To compare Equal Coverage and Proportional Representation, we report the Spearman correlation between these two measures. Specifically, we report the Spearman correlation between Equal Coverage value $EC(D,S)$ and Proportional Representation based on BARTScore of each summary generated by different LLMs. The results are in Tab. 9. We can observe that these two measures are not correlated for most datasets.
A.9 Lower and Upper bounds of the measure
@@
A.10 Summary-level Fairness under Different Distributions of Social attributes
To measure differences of the summary-level fairness measured by Equal Coverage under different distributions, we use maximum differences of equal coverage values $EC(G_{k})$ (Sec. 3.2) on sets dominated by different social attribute values $G_{k}$. The results are shown in Tab. 11. From the table, we can observe that summary-level fairness are relatively stable under different distributions of social attribute among input document set.
Ama. MIT. Art. Bia. Sem. New. Sta.
GPT-3.5 0.023 0.004 0.005 0.004 0.022
GPT-4 0.023 0.009 0.015 0.007 0.018
@@ A.10 Summary-level Fairness under Different Distributions of Social attributes
Claude3-son. 0.016 0.003 0.060 0.004 0.028
Table 11: Differences of Equal Coverage under different distributions of social attribute values in input document sets. Statistically significant ($p<0.05$) differences based on Bootstrapping are underlined. We observe significant differences in Equal Coverage on some datasets.

### v2

@@ Abstract
Fairness in multi-document summarization (MDS) measures whether a system can generate a summary fairly representing information from documents with different social attribute values. Fairness in MDS is crucial since a fair summary can offer readers a comprehensive view. Previous works focus on quantifying summary-level fairness using Proportional Representation, a fairness measure based on Statistical Parity. However, Proportional Representation does not consider redundancy in input documents and overlooks corpus-level unfairness. In this work, we propose a new summary-level fairness measure, Equal Coverage, which is based on coverage of documents with different social attribute values and considers the redundancy within documents. To detect the corpus-level unfairness, we propose a new corpus-level measure, Coverage Parity. Our human evaluations show that our measures align more with our definition of fairness. Using our measures, we evaluate the fairness of thirteen different LLMs. We find that Claude3-sonnet is the fairest among all evaluated LLMs. We also find that almost all LLMs overrepresent different social attribute values. The code is available at https://github.com/leehaoyuan/coverage_fairness.
@@ 3.2 Equal Coverage
However, two issues arise with summaries’ complex sentence structures. First, a summary sentence can combine information from several documents, making it difficult to attribute the sentence to any single document. Second, sentence lengths vary significantly based on social attribute values. The length difference can skew the coverage probability for different social attribute values $p(d,s|a=k)$. To address these issues, Equal Coverage splits the summary sentences $s\in S$ into multiple simpler sentences by prompting GPT-3.5 motivated by Bhaskar et al. (2023); Min et al. (2023). For example, compound sentences are split into simple sentences that describe a single fact. To reduce redundancy among split sentences, we filter out sentences with similar meanings based on entailment between split sentences. We denote the $j$-th summary sentence after split as $s_{j}$. Further details are in App. A.1.
Equal Coverage estimates the coverage probability for different social attribute values $p(d,s|a=k)$ and for a document $p(d,s)$ based on the probability $p(d_{i},s_{j})$ that a document $d_{i}$ is covered by a summary sentence $s_{j}$. The probability $p(d_{j},s_{k})$ is estimated as the entailment probability that the document $d_{j}$ entails the summary sentence $s_{k}$ by a textual entailment model Williams et al. (2018). Motivated by Laban et al. (2022), Equal Coverage divides documents into chunks of $M$ words. The $l$-th chunk of the document $d_{i}$ is denoted as $d_{i,l}$. The entailment model estimates the probability $p(d_{i},s_{j})$ as the maximum entailment probability $p(d_{i,l},s_{j})$ between the chunk $d_{i,l}$ and the summary sentence $s_{j}$:
@@ 4.2 Implementation Details
To measure the difficulty of obtaining fair summaries on different datasets, we estimate lower bounds, Lowergre, and upper bounds, Uppergre, for both fairness measures by greedily extracting sentences that minimize or maximize the measures. We also estimate random bounds, Random, by randomly extracting sentences from input document sets as summaries. Please note that Lowergre and Uppergre are not baselines but serve as empirical bounds for our measures. The details of these bounds are in App. A.8.
@@ 5.2 Summary-level Fairness Evaluation
Previous works by Zhang et al. (2023) and Lei et al. (2024) evaluate the fairness of LLMs using Proportional Representation. When evaluating the fairness on sentiments in the review domain, our results using Equal Coverage are consistent with these works in the finding that GPT-4 is fairer than GPT-3.5, and both are fairer than COOP. However, they find that Llama2-13b is the fairest, while our results show that Llama2-13b and Llama2-70b are comparably fair. When evaluating the fairness on political ideologies in the tweet domain, our results are consistent with these works on that smaller models are fairer for Llama2. However, they find that GPT-4 is fairer than GPT-3.5 while we find that GPT-3.5 is fairer than GPT-4. We also show that GPT-4 is fairer than Llama2-7b, which contrasts the previous works. We additionally report the results for Proportional Representation on our datasets in App. A.9 and also find much inconsistency between Equal Coverage and Proportional Representation. As shown in Sec. 1, Equal Coverage considers redundancy common in MDS, suggesting better reflects fairness in MDS.
@@ 5.4 Fainess under Different Distributions of Social Attributes
Ama. MIT. Art. Bia. Sem. New. Sta.
GPT-3.5 0.023 0.004 0.005 0.004 0.022
GPT-4 0.023 0.009 0.015 0.007 0.018
[…]
Llama2-13b 0.012 0.006 0.010 0.004 0.004
Llama2-70b 0.020 0.001 0.009 0.003 0.010
[…]
Llama3.1-70b 0.019 0.007 0.010 0.003 0.020
[…]
Mixtral-8x7b 0.010 0.003 0.003 0.004 0.012
[…]
Gemma2-27b 0.019 0.003 0.003 0.001 0.004
[…]
Claude3-son. 0.016 0.003 0.060 0.004 0.028
Table 4: Differences of Equal Coverage under different distributions of social attribute values in input document sets. Statistically significant ($p<0.05$) differences based on Bootstrapping are underlined. We observe significant differences in summary-level fairness for most LLMs on the Amazon dataset.
@@
We perform experiments to evaluate whether the fairness of LLMs changes under different distributions of social attribute values in input document sets. For this, we divide all samples $G$ into $K$ non-overlapping sets: $\{G_{1},...G_{K}\}$ based on distributions of social attribute values. Each set $G_{k}$ includes samples where most documents $d\in D$ have a social attribute value of $k$. We denote the set $G_{k}$ as the set dominated by social attribute value $k$. To measure differences of the summary-level fairness measured by Equal Coverage under different distributions, we use maximum differences of equal coverage values $EC(G_{k})$ (Sec. 3.2) on sets dominated by different social attribute values $G_{k}$. The results are shown in Tab. 4. From the table, we can observe that summary-level fairness changes much on the Amazon dataset but is relatively stable on the other dataset under different distributions of social attribute values.
To measure differences of the corpus-level fairness measured by Coverage Parity under different distributions, we use maximum differences of average coverage probability differences for the same social attribute value $\mathbb{E}(C_{j})$ (Sec. 3.3) on different sets $G_{k}$. The results are in Table 5.
@@ 7 Limitations
8 Acknowledgment
The authors appreciate the helpful feedback from Somnath Basu Roy Chowdhury, Anvesh Rao Vijjini, and Anneliese Brei in the early phase of this work. This work was supported in part by NSF grant DRL-2112635 and 2338418.
@@ A.1 Split and Rephrase Summary Sentences
Motivated by Bhaskar et al. (2023); Min et al. (2023), we split all summary sentences into simple sentences. The goal of this step is to ensure that each summary sentence after the split only discusses a single fact. Specifically, compound sentences are split into simple sentences, while sentences with compound subjects or objects are split into sentences with simple subjects or objects. However, such a split might generate redundant sentences. To remove redundancy, we construct an entailment graph among split summary sentences. There will be an edge between two summary sentences if two summary sentences entail each other with an entailment probability higher than $0.95$ according to the entailment model. For each connected component of the graph, we only keep one sentence that entails most other sentences in the connected components. After that, summary sentences are rephrased to remove the reported speech, like ‘documents say what’. For splitting and rephrasing, we prompt GPT-3.5 with demonstrations. We show the example prompt and the example summary sentences after splitting and rephrasing in Fig. 2.
@@ A.4 Datasets
Mohammad et al. (2016) consists of tweets with labels of stance toward a target phrase such as Climate Change or Hillary Clinton. We cluster tweets about the same short phrase based on their TFIDF similarity into clusters. We then divide these clusters into input document sets of 20 tweets about the same target phrase.
@@ A.7 Fair Summarization Prompts
A.8 Bounds of the Fairness Measures
@@
For both fairness measures, we also estimate a random bound, Random, by randomly sampling sentences from input document sets until reaching the token limit.
A.9 Comparison between Proportional Representation and Equal Coverage
To compare Equal Coverage and Proportional Representation, we report the Proportional Representation on our datasets. Specifically, we use BARTScore to estimate the distribution of social attributes among summaries following Zhang et al. (2023). We then report the average distribution differences between summaries and input document sets in Tab. 11.
Ama. MIT. Art. Bia. Sem. New. Sta. Ove.
GPT-3.5 0.204 0.178 0.243 0.201 0.304 0.107
GPT-4 0.209 0.176 0.259 0.215 0.335 0.323
@@ A.9 Comparison between Proportional Representation and Equal Coverage
Claude3-sonnet 0.219 0.190 0.261 0.186 0.317 0.225
Table 11: Summary-level fairness measured by Proportional Representation and overall scores on different datasets. A lower value indicates a fairer system. Bold indicates the fairest system.
Compared with the results based on Equal Coverage (Tab. 2), the fairness measured by Proportional Representation is very different. While Equal Coverage finds that larger models are generally fairer except for the family of Llama 3.1-8b, Proportional Representation only finds the same pattern for Claude3. Besides, Equal Coverage finds that Gemma2-27b is the fairest while Proportional Representation finds that GPT-3.5 is the fairest. We additionally report the Spearman correlation between Equal Coverage value $EC(D,S)$ and Proportional Representation in Tab. 12. From the table, we can observe that these two measures are not correlated for most datasets.
Amazon MITweet Article Bias SemEval News Stance
GPT-3.5 -0.025 -0.137 0.042 0.097 0.046
Llama2-70b 0.015 -0.091 0.172 0.078 -0.011
Mixtral-8x7b -0.006 -0.085 0.320 -0.072 0.083
Claude3-haiku -0.066 -0.026 0.071 -0.006 0.063